4. If the roots of the cubic az
+ bz + cz + d = 0 form an arithmetic progression α − β,
2
2
α, α + β, prove that (2b
− 9ac)b + 27a
d = 0.
1
Expert's answer
2020-05-24T16:20:31-0400

We are given that "az^3+bz^2+cz+d=0"  has three roots in arithmetic progression

"\alpha-\beta,\alpha,\alpha+\beta."

We know that


"z_1+z_2+z_3=-{b \over a}""z_1z_2z_3=-{d \over a}""z_1z_2+z_2z_3+z_1z_3={c \over a}"

Then


"\alpha-\beta+\alpha+\alpha+\beta=-{b \over a}"

"(\alpha-\beta)(\alpha)(\alpha+\beta)=-{d \over a}"

"(\alpha-\beta)(\alpha)+(\alpha)(\alpha+\beta)+(\alpha-\beta)(\alpha+\beta)={c \over a}"

We have


"\alpha=-{b \over 3a}"

"\alpha(\alpha^2-\beta^2)=-{d \over a}"

"3\alpha^2-\beta^2={c \over a}"

"\beta^2=3\alpha^2-{c \over a}={b^2 \over 3a^2}-{c \over a}"

Then


"-{b \over 3a}({b^2 \over 9a^2}-{b^2 \over 3a^2}+{c \over a})=-{d \over a}"

"{b(2b^2-9ac )\over 9a^2}=-3d"

"(2b^2-9ac )b+27a^2d=0"